OpenStudy (anonymous):

who can answer ma question plze: the mean weight of 8 boys is 55kg and the mean weight of a group of girls is 52kg . the mean weight of all the children is 53.2kg. how many girls are there?????

OpenStudy (anonymous):

12

OpenStudy (ash2326):

total weight of boys=55*8=440 let the no. of girls be x their total weight =x*52 now mean weight of all is 53.2 so (440+52x)/(x+8)=53.2 440+52x=53.2x+425.6 14.4=1.2x x=12 so no. of girls are 12

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

Thanks to Ash2326 for ur kind help

OpenStudy (anonymous):

wht did u wan tto "privatly" chat about?

OpenStudy (anonymous):

just to have more private fun!!!!!!!!!!!1

OpenStudy (anonymous):

hehehe *removes mind from gutter* wellll we can just tlk here w/ out ppl intudin then u can delete it

OpenStudy (anonymous):

ok this seem to be better place

OpenStudy (anonymous):

uhuh...soo tell me bout ur self (btw i still dont believe u to b 20)

OpenStudy (anonymous):

ok iam not that much old

OpenStudy (anonymous):

may i give u ma email now

OpenStudy (anonymous):

apparently u cant even form sentances properly no offense

OpenStudy (anonymous):

what do u mean by that?

OpenStudy (anonymous):

u either type sooo fast u cant keep up or u are lieing by u age to the 10th degree

OpenStudy (anonymous):

no way

OpenStudy (anonymous):

do u speak english as ur first language

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

that is gud

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

r u their lets

OpenStudy (anonymous):

r u there.......their is lik "tht is thier dog.." and yeah we had a bathroom break

OpenStudy (anonymous):

hey u there?

OpenStudy (anonymous):

yap am here

OpenStudy (anonymous):

be honest how old r u...no lie pls....

OpenStudy (anonymous):

17

OpenStudy (anonymous):

and u plze

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

am here and waiting u!!!!!!!!!!!!11

OpenStudy (anonymous):

16

OpenStudy (anonymous):

and what about ur nationality

OpenStudy (anonymous):

and believes

OpenStudy (anonymous):

american i live in NC wbu

OpenStudy (anonymous):

gud , and what do u believe

OpenStudy (anonymous):

i mean ur religion

OpenStudy (anonymous):

i am a baptist....

OpenStudy (anonymous):

is a baptist religion?

OpenStudy (anonymous):

i have never heard before

OpenStudy (anonymous):

yeah christian....wbu

OpenStudy (anonymous):

mmmmmmmmmmm

OpenStudy (anonymous):

? where do u live

OpenStudy (anonymous):

east Africa

OpenStudy (anonymous):

wow!

OpenStudy (anonymous):

iam going home now so c u

OpenStudy (anonymous):

so u can add up

OpenStudy (anonymous):

is that ok

OpenStudy (anonymous):

mmmkay i guess....dont 4get to delete tht

OpenStudy (anonymous):

ok and many thanks to u

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

wht is kkkk is all abt

OpenStudy (anonymous):

its ok 2x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

and bye bye